OpenStudy (anonymous):

Let f(x)=x^3−6x^2+4x−7. Find the open intervals on which f is concave up (down). Then determine the x-coordinates of all inflection points of f. 1. f is concave up on the intervals 2. f is concave down on the intervals 3. The inflection points occur at x =

OpenStudy (anonymous):

\[ f(x)=x^3−6x^2+4x−7\] did you take the derivative?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then you have to take the derivative again

OpenStudy (anonymous):

by which i mean you need the second derivative it will be a line (degree 1)

OpenStudy (anonymous):

what do you get for \(f''(x)\) ?

OpenStudy (anonymous):

okay i took the second derivative how im suppose to find the concavity

OpenStudy (anonymous):

if the second derivative is positive, then the function is "concave up"

OpenStudy (anonymous):

and if the second derivative is negative, the function is concave down the inflexion point is where the second derivative is 0

OpenStudy (anonymous):

but i think i have to plug in critical points how do i solve for those

OpenStudy (anonymous):

nah no need for critical points it just asks "concave up" "concave down" "inflexion"

OpenStudy (anonymous):

what did you get for the second derivative?

OpenStudy (anonymous):

6x-12 .... but then how would i know hat my intervals are

OpenStudy (anonymous):

or actually "inflection" i cant spell too good

OpenStudy (anonymous):

if you want to know over what interval the function is concave up, solve \[6x-12>0\] which you can almost do in your head

OpenStudy (anonymous):

so 2 would be that its positive and that means its concave up

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

you are looking for an interval, not a number, an interval over which the function is concave up the second derivative is \(6x-12\) and the function will be concave up if \(6x-12>0\) i.e. it is concave up over the interval for which \(6x-12>0\) if \(x>2\) then \(6x-12>0\) and so the function is concave up on the interval \[(2,\infty)\]

OpenStudy (anonymous):

the fact that 2 is positive has nothing to do with the function being concave up, it is just the endpoint of the interval over which \(6x-12>0\)

OpenStudy (anonymous):

and of course if \(6x-12<0\) that means \(x<2\) so the function is concave down on \[(-\infty,2)\]

OpenStudy (anonymous):

if you think about what a cubic function with positive leading coefficient looks like, this makes perfect sense 

OpenStudy (anonymous):

my point of inflection would be at 2 right

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw